OpenStudy (anonymous):

List the first four terms of a geometric sequence with t1 = 4 and tn = -3tn-1. im lost...again

OpenStudy (anonymous):

@ospreytriple @Mehek14

OpenStudy (campbell_st):

so for the 2nd term substitute n = 2 t2 = -3 x t1 so t2 = -3 x 4 t2 = -12 and then repeat the process for n =3 and n = 4 hope it helps

OpenStudy (anonymous):

For \(t_2\), put 2 in for \(n\) in the given formula, i.e.\[t_2 = -3\times t_{2-1} = -3\times t_1\]and you know what \(t_1\) is.

OpenStudy (anonymous):

i got -12, is there more i have to do?

OpenStudy (anonymous):

Correct. Now you have \(t_1\) and \(t_2\). The question asks for the first four terms, so you need to repeat what you did above and calculate \(t_3\) and \(t_4\).

OpenStudy (anonymous):

so for t3 i do t3=-12 x t3-1?

OpenStudy (anonymous):

Not quite. The formula is\[t_n = -3 t_{n-1}\]The -3 is a constant, it never changes. In fact, when you understand the pattern, you'll see that -3 is the common ratio.

OpenStudy (anonymous):

is t3=-3t3-1 t3=-3t2 t3=-3(-12) t3=36 like that?

OpenStudy (anonymous):

Excellent! Now, what's \(t_4\) ?

OpenStudy (anonymous):

t4=-3t4-1 -3t3 -3(36) -108 like that?

OpenStudy (anonymous):

Congrats! Problem finished.

OpenStudy (anonymous):

thank you!. im glade you someone who shows me then just tells me the anwser

OpenStudy (anonymous):

I agree. You're welcome.

OpenStudy (anonymous):

just woundering, but are you good with ellipses and hyperbola?

OpenStudy (anonymous):

okay i have another question...ill start another one